OpenStudy (anonymous):

Solve the equation explicitly for y and and differentiate to get y' in terms of x. (x^(1/2))+(y^(1/2))=1

OpenStudy (anonymous):

\[\sqrt{x}+\sqrt{y}=1\]

OpenStudy (anonymous):

So this is what I thought they are trying to get at: \[\sqrt{y}=1-\sqrt{x}\] square root both sides to take other the square root on y. \[[\sqrt{y}=1-\sqrt{x}]^{2}\] \[y=1+x\] \[y'=1\] This is wrong.

OpenStudy (freckles):

\[(1-\sqrt{x})^2 \neq 1+x\]

OpenStudy (anonymous):

oh it doesnt?

OpenStudy (freckles):

\[(1-\sqrt{x})^2=(1-\sqrt{x})(1-\sqrt{x})=1-\sqrt{x}-\sqrt{x}+x=1-2 \sqrt{x}+\sqrt{x}\] but I would just leave it as (1-sqrt(x))^2 and use chain rule

OpenStudy (anonymous):

ok one second let me change that up and see what I get now.

OpenStudy (freckles):

oops I made a type-0

OpenStudy (freckles):

that last term was suppose to be x

OpenStudy (anonymous):

yea the last sqrt x should be just x

OpenStudy (freckles):

\[1-2 \sqrt{x}+x\]

OpenStudy (anonymous):

I got \[\frac{ 1-\sqrt{x} }{ \sqrt{x} }\]

OpenStudy (anonymous):

is that what you got as well?

OpenStudy (freckles):

hmmm I got the opposite of that

OpenStudy (freckles):

which just means our answers are off by a constant multiple of 1

OpenStudy (anonymous):

oh wait I forgot the negative in my answer

OpenStudy (freckles):

\[\sqrt{x}+\sqrt{y}=1 \\ \sqrt{y}=1-\sqrt{x} \\ y=(1-\sqrt{x})^2 \\ y'=2(1-\sqrt{x}) \cdot (0- \frac{1}{2 \sqrt{x}}) \\ y'=\frac{-2(1-\sqrt{x})}{2 \sqrt{x}} =\frac{-(1-\sqrt{x})}{\sqrt{x}}\]

OpenStudy (anonymous):

yup now I got that as well

OpenStudy (freckles):

cool stuff

OpenStudy (anonymous):

alright thank you again :D

OpenStudy (freckles):

np you are welcome again